OpenStudy (anonymous):

To fit between two windows, the width of a bookshelf must be no greater than feet. Mrs. Aguilar purchases a bookshelf that is 77 inches wide. Which statement describes the relationship between the width of the bookshelf and the distance between the windows?

OpenStudy (anonymous):

To fit between two windows, the width of a bookshelf must be no greater than 6.5 feet. Mrs. Aguilar purchases a bookshelf that is 77 inches wide. Which statement describes the relationship between the width of the bookshelf and the distance between the windows? You left out the 6.5 feet that's why nobody answered your question, next time check your question and make sure you have all of the information so people actually know what you are talking about. The width between the two windows is 6.5 feet, which is 78 inches. If Mrs. Aguilar purchased a bookshelf that's 77 inches wide, her bookshelf will fit between the two windows.

OpenStudy (anonymous):

self width (s) must be ≤ to distance (d) between windows. d is known 6.5 feet. Mrs. Aguilar can buy any shelf whos width makes the statement below true; s≤d d =6.5 feet in this specific case for the self measuring 78 inches you will need to convert ur units and then compare; \[(78 inch) \frac{ 1 foot }{ 12 inch}\le 6.5 feet\]

OpenStudy (anonymous):

correction \[77 inch \frac{ 1 foot }{12inch }\le 6.5 feet\]